THE MAGELLANIC CLOUDS NEWSLETTER

An electronic publication dedicated to the Magellanic Clouds, and astrophysical phenomena therein

No. 104 — 2 April 2010

http://www.astro.keele.ac.uk/MCnews

Editor: Jacco van Loon

Editorial

Dear Colleagues, It is my pleasure to present you the 104^{th} issue of the Magellanic Clouds Newsletter. Marvel at the MAD images of the heart of the Tarantula Nebula’s nest of massive stars, or be amazed by the three (sic!) papers on SN 1987 A alone (plus one about its surrounding field) and the two papers on supernova remnants and their progenitors by Maoz & Badenes (and Draine). The PDF version of this newsletter contains a two-page advertisement for a special issue of the Royal Society’s Philosophical Transactions, on the topic of star clusters. There is an attractive discount being offered on buying this book. This issue does not feature a cover picture as no suggestions were received. They remain very welcome for future issues. The next issue is planned to be distributed on the 1^{st} of June 2010.

Editorially Yours, Jacco van Loon

Refereed Journal Papers

VLT-MAD observations of the core of 30 Doradus

M.A. Campbell^{1}, C.J. Evans^{2,1}, A.D. Mackey^{1}, M. Gieles^{3}, J. Alves^{4}, J. Ascenso^{5}, N. Bastian^{6} and A.J. Longmore^{2}

1Institute for Astronomy, The University of Edinburgh, Royal Observatory, Blackford Hill, Edinburgh EH9 3HJ, UK 2UK Astronomy Technology Centre, Royal Observatory, Blackford Hill, Edinburgh, EH9 3HJ, UK 3European Southern Observatory, Casilla 19001, Santiago 19, Chile 4Calar Alto Observatory-Centro Astronómico Hispano-Alemán, C/ Jesús Durbán Remón 2-2, 04004 Almeria, Spain 5Harvard-Smithsonian Center for Astrophysics, 60 Garden Street, Cambridge, MA 02138, USA 6Institute of Astronomy, University of Cambridge, Madingley Road, Cambridge, CB3 0HA, UK

We present H- and K_{s}-band imaging of three fields at the centre of 30 Doradus in the Large Magellanic Cloud, obtained as part of the Science Demonstration programme with the Multi-conjugate Adaptive optics Demonstrator (MAD) at the Very Large Telescope. Strehl ratios of 15–30% were achieved in the K_{s}-band, yielding near-infrared images of this dense and complex region at unprecedented angular resolution at these wavelengths. The MAD data are used to construct a near-infrared luminosity profile for R 136, the cluster at the core of 30 Dor. Using cluster profiles of the form used by Elson et al., we find the surface brightness can be fit by a relatively shallow power-law function (γ ∼ 1.5–1.7) over the full extent of the MAD data, which extends to a radius of ∼ 40′′ (∼ 10 pc). We do not see compelling evidence for a break in the luminosity profile as seen in optical data in the literature, arguing that cluster asymmetries are the dominant source, although extinction effects and stars from nearby triggered star-formation likely also contribute. These results highlight the need to consider cluster asymmetries and multiple spatial components in interpretation of the luminosity profiles of distant unresolved clusters. We also investigate seven candidate young stellar objects reported by Gruendl & Chu from Spitzer observations, six of which have apparent counterparts in the MAD images. The most interesting of these (GC09 053839.24−690552.3) appears related to a striking bow-shock–like feature, orientated away from both R 136 and the Wolf-Rayet star Brey 75, at distances of 19.5′′ and 8′′ (4.7 and 1.9 pc in projection), respectively.

Accepted for publication in MNRAS Available from arXiv:1002.0288

New tests for disruption mechanisms of star clusters: The Large and Small Magellanic Clouds

Rupali Chandar^{1}, S. Michael Fall^{2} and Bradley C. Whitmore^{2}

1Department of Physics & Astronomy, University of Toledo, Toledo, OH 43606, USA 2Space Telescope Science Institute, 3700 San Martin Dr., Baltimore, MD 21218, USA

We compare the observed bivariate distribution of masses (M) and ages (t) of star clusters in the Large Magellanic Cloud (LMC) with the predicted distributions g(M, t) from three idealized models for the disruption of star clusters:

(1) sudden mass-dependent disruption; (2) gradual mass-dependent disruption; and (3) gradual mass-independent disruption. The model with mass-independent disruption provides a good, first-order description of these cluster populations, with g(M, t) ∝ M^{β}t^{γ}, β = −1.8 ± 0.2 and γ = −0.8 ± 0.2, at least for clusters with ages t < 10^{9} yr and masses M < few ×10^{3} M_{⊙} (more specifically, t < 10^{7}(M/10^{2} M_{⊙})^{1.3} yr). This model predicts that the clusters should have a power-law luminosity function, dN/dL ∝ L^{−1.8}, in agreement with the observations. The first two models, on the other hand, fare poorly when describing the observations, refuting previous claims that mass-dependent disruption of star clusters is observed in the LMC over the studied M–t domain. Clusters in the SMC can be described by the same g(M, t) distribution as for the LMC, but with smaller samples and hence larger uncertainties. The successful g(M, t) model for clusters in the Magellanic Clouds is virtually the same as the one for clusters in the merging Antennae galaxies, but extends the domain of validity to lower masses and to older ages. This indicates that the dominant disruption processes are similar in these very different galaxies over at least t < 10^{8} yr and possibly t < 10^{9} yr. The mass functions for young clusters in the LMC are power-laws, while that for ancient globular clusters is peaked. Ancient clusters in the LMC have a higher mean density than clusters with ages between t = 10^{8}–10^{9} yr. We show that the observed shapes of the mass functions are consistent with expectations from the simple evaporation model presented by McLaughlin & Fall.

Accepted for publication in Astrophysical Journal Available from arXiv:1002.0779

The column density towards LMC X-1

Manfred Hanke^{1}, J. Wilms^{1}, Michael A. Nowak^{2}, Laura Barragán^{1} and Norbert S. Schulz^{2}

1Dr. Karl Remeis-Observatory / ECAP, University of Erlangen-Nuremberg, Sternwartstr. 7, 96049 Bamberg, Germany 2MIT Kavli Institute for Astrophysics and Space Research, NE80-6077, 77 Mass. Ave., Cambridge, MA 02139, USA

We measure the neutral absorption towards the black hole X-ray binary system LMC X-1 from six archival soft X-ray spectra obtained with the gratings and/or CCD detectors on Chandra, XMM-Newton, and Swift. Four spectral models for the soft continuum are investigated. While the powerlaw model may overestimate N_{H} considerably, the others give consistent results. Taking the lower metalicity of the Large Magellanic Cloud into account, we find equivalent hydrogen column densities of N_{H} = (1.0 − 1.3) × 10^{22} cm^{−2}, with a systematic dependence on the orbital phase. This variation in the neutral absorption can nearly explain the orbital modulation of the soft X-ray flux recently detected with the All Sky Monitor (ASM) on the Rossi X-ray Timing Explorer (RXTE).

Published in Astronomy and Astrophysics, 509, L8 (2010) Available from arXiv:0912.3719

High Resolution X-Ray Spectroscopy of SN 1987 A: Monitoring with XMM-Newton

Richard Sturm^{1}, Frank Haberl^{1}, Bernd Aschenbach^{1} and Günther Hasinger^{1,2}

1Max-Planck-Institut für extraterrestrische Physik, Giessenbachstraße, 85748 Garching, Germany 2Max-Planck-Institut für Plasmaphysik, Boltzmannstraße 2, 85748 Garching, Germany

We report the results of our XMM-Newton monitoring of SN 1987 A. The ongoing propagation of the supernova blast wave through the inner circumstellar ring caused a drastic increase in X-ray luminosity during the last years, enabling detailed high resolution X-ray spectroscopy with the Reflection Grating Spectrometer. The observations can be used to follow the detailed evolution of the arising supernova remnant. The fluxes and broadening of the numerous emission lines seen in the dispersed spectra provide information on the evolution of the X-ray emitting plasma and its dynamics. These were analyzed in combination with the EPIC-pn spectra, which allow a precise determination of the higher temperature plasma. We modeled individual emission lines and fitted plasma emission models. Especially from the observations between 2003 and 2007 we can see a significant evolution of the plasma parameters and a deceleration of the radial velocity of the lower temperature plasma regions. We found an indication (3σ-level) of an iron K feature in the co-added EPIC-pn spectra. The comparison with Chandra grating observations in 2004 yields a clear temporal coherence of the spectral evolution and the sudden deceleration of the expansion velocity seen in X-ray images ∼6100 days after the explosion.

Accepted for publication in A&A Available from arXiv:1002.1865

Observational constraints on the modified gravity model (MOG) proposed by Moffat: using the Magellanic System

Hosein Haghi^{1} and Sohrab Rahvar^{2,3}

1Institute for Advanced Studies in Basic Sciences (IASBS), P.O. Box 45195-1159, Zanjan, Iran 2Department of Physics, Sharif University of Technology, P.O. Box 11365-9161, Tehran, Iran 3School of Astronomy and Astrophysics, Institute for Research in Fundamental Sciences, (IPM), P.O. Box 19395-5531, Tehran, Iran

A simple model for the dynamics of the Magellanic Stream (MS), in the framework of modified gravity models is investigated. We assume that the galaxy is made up of baryonic matter out of context of dark matter scenario. The model we used here is named Modified Gravity (MOG) proposed by Moffat (2005). In order to examine the compatibility of the overall properties of the MS under the MOG theory, the observational radial velocity profile of the MS is compared with the numerical results using the 2 fit method. In order to obtain the best model parameters, a maximum likelihood analysis is performed. We also compare the results of this model with the Cold Dark Matter (CDM) halo model and the other alternative gravity model that proposed by Bekenstein (2004), so called TeVeS. We show that by selecting the appropriate values for the free parameters, the MOG theory seems to be plausible to explain the dynamics of the MS as well as the CDM and the TeVeS models.

Accepted for publication in International Journal of Theoretical Physics Available from arXiv:1002.3136

Photometric determination of the mass accretion rates of pre-main sequence stars. I. Method and application to the SN 1987 A field

Guido De Marchi^{1}, Nino Panagia^{2,3,4} and Martino Romaniello^{5}

1European Space Agency, Space Science Department, Keplerlaan 1, 2200 AG Noordwijk, The Netherlands 2Space Telescope Science Institute, 3700 San Martin Drive, Baltimore, MD 21218, USA 3INAF-CT, Osservatorio Astrofisico di Catania, Via S. Sofia 78, 95123, Italy 4Supernova Limited, VGV #131, Northsound Road, Virgin Gorda, British Virgin Islands 5European Southern Observatory, Karl-Schwarzschild-Strasse 2, 85748 Garching, Germany

We have developed and successfully tested a new self-consistent method to reliably identify pre-main sequence (PMS) objects actively undergoing mass accretion in a resolved stellar population, regardless of their age. The method does not require spectroscopy and combines broad-band V and I photometry with narrow-band Hα imaging to: (1) identify all stars with excess Hα emission; (2) derive their Hα luminosity L(Hα); (3) estimate the Halpha emission equivalent width; (4) derive the accretion luminosity L_{acc} from L(Hα); and finally (5) obtain the mass accretion rate M_{acc} from L_{acc} and the stellar parameters (mass and radius). By selecting stars with photometric accuracy in Hα better than 15%, the statistical uncertainty on the derived M_{acc} is typically < 17% and is dictated by the precision of the Hα photometry. Systematic uncertainties, of up to a factor of 3 on the value of M_{acc}, are caused by our incomplete understanding of the physics of the accretion process and affect all determinations of the mass accretion rate, including those based on a spectroscopic Hα line analysis. As an application of our method, we study a field of 9.16 arcmin^{2} around SN 1987 A, using existing HST photometry. We identify as bona-fide PMS stars a total of 133 objects with a Hα excess above the 4-σ level and a median age of 13.5 Myr. Their median mass accretion rate of 2.6 × 10^{−8} M_{⊙} yr^{−1} is in excellent agreement with previous determinations based on the U-band excess of the stars in this field, and with the value measured for G-type PMS stars in the Milky Way. Their L_{acc} shows a strong dependence on their distance from a group of hot massive stars in the field and suggests that the UV radiation of the latter is rapidly eroding the circumstellar discs around PMS stars.

Accepted for publication in Astrophysical Journal Available from arXiv:1002.4864

Abundances and density structure of the inner circumstellar ring around SN 1987 A

Seppo Mattila^{1,2}, Peter Lundqvist^{2}, Per Gröningsson^{2}, Peter Meikle^{3}, Raylee Stathakis^{4}, Claes Fransson^{2,5} and Russell Cannon^{4}

1Tuorla Observatory, Finland 2Stockholm Observatory, Sweden 3Imperial College London, UK 4AAO, Australia 5Oskar Klein Centre, Sweden

We present optical spectroscopic data of the inner circumstellar ring around SN 1987 A from the Anglo-Australian Telescope (AAT) and the Very Large Telescope (VLT) between ∼ 1400 and ∼ 5000 days post-explosion. We also assembled the available optical and near-infrared line fluxes from the literature between ∼ 300 and ∼ 2000 days. These line light curves were fitted with a photoionization model to determine the density structure and the elemental abundances for the inner ring. We found densities ranging from 1 × 10^{3} to 3 × 10^{4} atoms cm^{−3} and a total mass of the ionized gas of 0.058 M_{⊙} within the inner ring. Abundances inferred from the optical and near-infrared data were also complemented with estimates of Lundqvist & Fransson (1996) based on ultraviolet lines. This way we found an He/H-ratio (by number of atoms) of 0.17 ± 0.06, N/O-ratio of 1.5 ± 0.7, and the total (C+N+O)/(H+He) abundance about 1.6 times its LMC value or roughly 0.6 times the most recent solar value. An iron abundance of 0.20±0.11 times solar was found which is within the range of the estimates for the LMC. We also present late time (∼ 5000–7500 days) line light curves of [O iii], [Ne iii], [Ne iv], [Ar iii], [Ar iv], and [Fe vii] from observations with the VLT. We compared these with model fluxes and found that an additional 10^{2} atoms cm^{−3} component was required to explain the data of the highest ionization lines. Such low density gas is expected in the H ii-region interior to the inner ring which likely extends also to larger radii at higher latitudes. At epochs later than ∼ 5000 days our models underproduce the emission of most of these lines as expected due to the contribution from the interaction of the SN ejecta with the ring.

Submitted to ApJ Available from arXiv:1002.4195

Probing the LMC age gap at intermediate cluster masses

Eduardo Balbinot^{1}, Basílio Santiago^{1}, Leandro Kerber^{2}, Beatriz Barbuy^{2} and Bruno Dias^{2}

1Universidade Federal do Rio Grande do Sul, Porto Alegre, RS, Brazil 2Intituto de Astronomia de Geociências, USP, São Paulo, SP, Brazil

The LMC has a rich star cluster system spanning a wide range of ages and masses. One striking feature of the LMC cluster system is the existence of an age gap between 3–10 Gyrs. But this feature is not as clearly seen among field stars. Three LMC fields containing relatively poor and sparse clusters whose integrated colours are consistent with those of intermediate age simple stellar populations have been imaged in BV I with the Optical Imager (SOI) at the Southern Telescope for Astrophysical Research (SOAR). A total of 6 clusters, 5 of them with estimated initial masses M < 10^{4} M_{⊙}, were studied in these fields. Photometry was performed and Colour–Magnitude Diagrams (CMD) were built using standard point spread function fitting methods. The faintest stars measured reach V ∼ 23. The CMD was cleaned from field contamination by making use of the three-dimensional colour and magnitude space available in order to select stars in excess relative to the field. A statistical CMD comparison method was developed for this purpose. The subtraction method has proven to be successful, yielding cleaned CMDs consistent with a simple stellar population. The intermediate age candidates were found to be the oldest in our sample, with ages between 1–2 Gyrs. The remaining clusters found in the SOAR/SOI have ages ranging from 100 to 200 Myrs. Our analysis has conclusively shown that none of the relatively low-mass clusters studied by us belongs to the LMC age-gap.

Accepted for publication in MNRAS Available from arXiv:1002.0805

A new population of Planetary Nebulae discovered in the Large Magellanic Cloud. III: The luminosity function

Warren Reid^{1} and Quentin Parker^{1,2}

1Macquarie University, Sydney, Australia 2Anglo-Australian Observatory, Australia

Our previous identification and spectroscopic confirmation of 431 faint, new planetary nebulae in the central 25 deg^{2} region of the LMC permits us to now examine the shape of the LMC Planetary Nebula Luminosity Function (PNLF) through an unprecedented 10 magnitude range. The majority of our newly discovered and previously known PNe were observed using the 2dF, multi-object fibre spectroscopy system on the 3.9-m Anglo-Australian Telescope and the FLAMES multi-object spectrograph on the 8-m VLT. We present reliable [O iii] 5007 and Hβ flux estimates based on calibrations to well established PN fluxes from previous surveys and spectroscopic standard stars. The bright cutoff (M^{⋆}) of the PNLF is found by fitting a truncated exponential curve to the bright end of the PNLF over a 3.4 magnitude range. This cutoff is used to estimate a new distance modulus of 18.46 to the LMC, in close agreement with previous PNLF studies and the best estimates by other indicators. The bright end cutoff is robust to small samples of bright PNe since significantly increased PN samples do not change this fiducial. We then fit a truncated exponential curve directly to the bright end of the function over a 6 magnitude range and test the curve’s ability to indicate the position of M^{⋆}. Because of the significant increase in the number of LMC PN, the shape of the PNLF is now examined in greater detail than has previously been possible. Newly discovered features include a small increase in the number of PNe over the brightest 4 magnitudes followed by a steep rise over 2 magnitudes, a peak at 6 magnitudes below the bright cutoff and an almost linear drop-off to the faint end. Dips at the bright end of the PNLF are examined in relation to the overall shape of the PNLF and the exponential increase in the number of PNe. Through cumulative functions, the new LMC PNLF is compared to those from the SMC and a new deep local Galactic sample revealing the effects of incompleteness. The new [O iii] 5007 LMC PNLF is then compared to our new Hβ LMC PNLF using calibrated and measured fluxes for the same objects, revealing the effects of metallicity on the [O iii] 5007 line.

Accepted for publication in MNRAS Available from arXiv:1002.3410

A binary merger model for the formation of the Supernova 1987 A triple-ring nebula

Thomas Morris^{1,2} and Philipp Podsiadlowski^{1}

1Oxford University, UK 2Max-Planck Institute for Astrophysics, Garching bei München, Germany

We examine a binary merger model for the formation of the mysterious triple-ring nebula surrounding Supernova 1987 A, which still has not been convincingly explained in detailed hydrodynamical calculations. During the merger of 15 and 5-M_{⊙} binary systems, mass is ejected primarily at mid-latitudes for a sufficiently evolved primary, as demonstrated by Morris & Podsiadlowski (2005). This material is swept up by the fast wind of the central star during its post-merger blue supergiant phase, leading to a density contrast of ∼ 150 in the outer rings at the time of the supernova. The equatorial ring probably formed later when the star contracted to become a blue supergiant. The asymmetry between the northern and southern outer rings can be explained by a 10 per cent asymmetry during the merger, perhaps due to a pulsational instability in the common envelope. We present a parameter study from which we determine a mass-loss rate in the blue supergiant wind in the range 1.5 − 3 × 10^{−7} M_{⊙} yr^{−1} in agreement with previous estimates. The morphology of the best model is consistent with the well-known Hubble Space Telescope image at better than 5 per cent and is also in broad agreement with light-echo observations. The circumstellar environment on larger scales (up to 3 pc) is also investigated. We conclude with a brief discussion of the bipolar nebulae surrounding the Galactic stars, Sheridan 25, HD 168625 and η Carinae.

Published in MNRAS, 399, 515 (2009)

HST/WFPC2 observations of the LMC pulsar PSR B0540−69

Roberto Mignani^{1}, Angelica Sartori^{2}, Andrea de Luca^{2,3}, Bronek Rudak^{4,5}, Aga Slowikowska^{6,7}, Gottfried Kanbach^{8} and Patrizia Caraveo^{2}

1Mullard Space Science Laboratory, University College London, Holmbury St. Mary, Dorking, Surrey, RH5 6NT, UK 2INAF - Instituto di Astrofisica Spaziale e Fisica Cosmica di Milano,Via Bassini 15, I-20133 Milano, Italy 3Instituto Universitario di Studi Superiori Viale Lungo Ticino Sforza 56, Pavia, 27100, Italy 4Nicolaus Copernicus Astronomical Center, ul. Rabianska 8, 87100, Torun, Poland 5KAA UMK, Gagarina 11, 87-100 Torun, Poland. 6Institute of Astronomy, University of Zielona Gora, Lubuska 2, 65-265 Zielona Gora, Poland 7IESL, Foundation for Research and Technology, 71110 Heraklion, Crete, Greece 8Max-Planck Institut für Extraterrestrische Physik, Gießenbachstraße 1, 85741 Garching bei München, Germany

About 40 years after the identification of the optical counterpart of the Crab pulsar, only a few rotation-powered pulsars have been identified in the optical domain. Of these, only four have spin-down ages of < 10 000 years: the Crab pulsar, PSR B1509−58, PSR B0540−69 in the Large Magellanic Cloud (LMC), and the Vela pulsar. The study of the younger, and brighter, pulsars is important to understand the optical emission properties of isolated neutron stars. PSR B0540−69, the second brightest (V ∼ 22) optical pulsar, is obviously a very interesting target for these investigations. The aim of this work is threefold: constraining the pulsar proper motion and its velocity on the plane of the sky through optical astrometry, obtaining a more precise characterisation of the pulsar optical spectral energy distribution (SED) through a consistent set of multi-band, high-resolution, imaging photometry observations, measuring the pulsar optical phase-averaged linear polarisation, for which only a preliminary and uncertain measurement was obtained so far from ground-based observations. We performed high-resolution observations of PSR B0540−69 with the WFPC2 aboard the HST, in both direct imaging and polarimetry modes. From multi-epoch astrometry we set a 3-σ upper limit of 1 mas yr^{−1} on the pulsar proper motion, implying a transverse velocity < 250 km s^{−1} at the 50 kpc LMC distance. Moreover, we determined the pulsar absolute position with an unprecedented accuracy of 70 mas. From multi-band photometry we characterised the pulsar power-law spectrum and we derived the most accurate measurement of the spectral index (0.70 ± 0.07) which indicates a spectral turnover between the optical and X-ray bands. Finally, from polarimetry we obtained a new measurement of the pulsar phase-averaged polarisation degree (16 ± 4%),consistent with magnetosphere models depending on the actual intrinsic polarisation degree and depolarisation factor, and we found that the polarisation vector (22 ± 12^{∘} position angle) is possibly aligned with the semi-major axis of the pulsar-wind nebula and with the apparent proper motion direction of its bright emission knot.

Accepted for publication in Astronomy and Astrophysics Available from arXiv:1003.0786

The supernova rate and delay times in the Magellanic Clouds

Dan Maoz^{1} and Carles Badenes^{1,2}

1Tel-Aviv University, Tel-Aviv, Israel 2Weizmann Institute, Rehovot, Israel

We use the supernova remnants (SNRs) in the two Magellanic Clouds (MCs) as a supernova (SN) survey, “conducted” over tens of kyr, from which we derive the current SN rate, and the SN delay time distribution (DTD), i.e., the SN rate vs. time that would follow a hypothetical brief burst of a star formation. In a companion paper (Badenes, Maoz, & Draine 2010) we have compiled a list of 77 SNRs in the MCS, and argued that it is a largely complete record of the SNRs that are now in the Sedov phase of their expansions. We recover the SN DTD by comparing, on one hand, the numbers of SNRs observed in small individual “cells” in these galaxies to, on the other hand, the star-formation histories of each cell, as calculated from resolved stellar populations by Harris & Zaritsky. We identify the visibility times of SNRs with the Sedov-phase lifetimes, which depend on the local ambient densities. The local densities are estimated from 21-cm emission, from an inverse Schmidt-Kennicutt law based on either Hα emission or on the starformation rate from the resolved stellar populations, and from combinations of these tracers. This is the first SN DTD that is based on resolved stellar populations. We detect a population of “prompt” type-Ia SNe (that explode within 330 Myr of star formation) at > 99% confidence level (c.l.). The best fit for the number of prompt-SNe per stellar mass formed is 2.7–11.0 × 10^{−3} M_{⊙}^{−1}, depending on the density tracer used. The 95% c.l. range for a “delayed” (from 330 Myr to a Hubble time) type-Ia component is < 1.6 × 10^{−13} SN yr^{−1} M_{⊙}^{−1}, consistent with rate measurements in old populations. The current total (core-collapse+Ia) SN rate in the MCs is 2.5–4.6 SNe per millenium (68% c.l.+systematics), or 1.7–3.1 SNuM [SNe (100yr10^{10} M_{⊙})^{−1}], in agreement with the historical record and with rates measured in other dwarf irregulars. Conversely, assuming the SNRs are in free-expanion, rather than in their Sedov phase, would impose on the SNRs a maximum age of 6 kyr, and would imply a MC SN rate per unit mass that is 5 times higher than in any type of galaxy, and a low-mass limit for core-collapse progenitors in conflict with stellar evolution theory.

Submitted to MNRAS Available from arXiv:1003.3031

On the size distribution of supernova remnants in the Magellanic Clouds

Carles Badenes^{1,2}, Dan Maoz^{1} and Bruce Draine^{3}

1Tel-Aviv University, Tel-Aviv, Israel 2Weizmann Institute, Rehovot, Israel 3Princeton University, Princeton, NJ, USA

The physical sizes of supernova remnants (SNRs) in a number of nearby galaxies follow an approximately linear cumulative distribution, contrary to what is expected for decelerating shock fronts. This phenomenon has been variously attributed to observational selection effects, or to a majority of SNRs being in “free expansion”, with shocks propagating at a constant velocity into a tenuous ambient medium. We compile multi-wavelength observations of the 77 known SNRs in the Magellanic Clouds, and argue that they provide a largely complete record of the SNe that have exploded over the last ∼ 20 kyr, with most of them now in the adiabatic, Sedov phase of their expansions. The roughly linear cumulative distribution of sizes (roughly uniform in a differential distribution) can be understood to result from the combination of the deceleration during this phase, a transition to a radiation-loss-dominated phase at a radius that depends on the local gas density, and a probability distribution of densities in the interstellar medium varying approximately as ρ^{−1}. This explanation is supported by the observed powerlaw distributions, with index ∼ −1, of three independent tracers of density: neutral hydrogen column density, Hα surface brightness, and star-formation rate based on resolved stellar populations. In this picture, the observed cutoff at a radius of 30 pc in the SNR size distribution is due to a minimum in the mean ambient gas density in the regions where supernovae (SNe) explode. We show that M 33 has a SNR size distribution very similar to that of the Magellanic Clouds, suggesting these features, and their explanation, may be universal. In a companion paper (Maoz & Badenes 2010), we use our sample of SNRs as an effective “SN survey” to calculate the SN rate and delay time distribution in the Magellanic Clouds. The hypothesis that most SNRs are in free expansion, rather than in the Sedov phase of their evolution, would result in SN rates that are in strong conflict with independent measurements, and with basic stellar evolution theory.

Submitted to MNRAS Available from arXiv:1003.3030

RR Lyrae and Type II Cepheid variables adhere to a common distance relation

Daniel J. Majaess^{1,2}

1Saint Mary’s University, Halifax, Nova Scotia, Canada 2The Abbey Ridge Observatory, Stillwater Lake, Nova Scotia, Canada

Preliminary evidence is presented reaffirming that SX Phe, RR Lyrae, and Type II Cepheid variables may be characterized by a common Wesenheit period–magnitude relation, to first order (see also Matsunaga et al. 2006). Reliable distance estimates to RR Lyrae variables and Type II Cepheids are ascertained from a single V I-based reddening-free relation derived recently from OGLE photometry of LMC Type II Cepheids. Distances are computed to RR Lyrae (d ∼ 260 pc) and variables of its class in the galaxies IC 1613, M 33, Fornax dSph, LMC, SMC, and the globular clusters M 3, M 15, M 54, ω Cen, NGC 6441, and M 92. The results are consistent with literature estimates, and in the particular cases of the SMC, M 33, and IC 1613, the distances agree with those inferred from classical Cepheids to within the uncertainties: no corrections were applied to account for differences in metallicity. Moreover, no significant correlation was observed between the distances computed to RR Lyrae variables in ω Cen and their metallicity, despite a considerable spread in abundance across the sample. In sum, concerns regarding a sizeable metallicity effect are allayed when employing V I-based reddening-free Cepheid and RR Lyrae relations.

Accepted for publication in JAAVSO Available from http://www.aavso.org/publications/ejaavso/ej112.shtml